# General Flow as Foundation Affordance for Scalable Robot Learning

### v1

@@ VI Zero-shot Real World Manipulation
In this section, we address one of the most challenging tasks: stable zero-shot human-to-robot skill transfer in real-world scenarios, demonstrating the foundational capabilities of general flow prediction. Utilizing only a single prediction model paired with a straightforward heuristic policy derived from closed-loop prediction, we achieve an impressive 81% average success rate. This success spans categories including rigid, articulated, and soft bodies and covers 18 tasks across 6 distinct scenes. Deriving more effective policies through approaches like few-shot imitation learning [85] or reinforcement learning [4] is left for future work.
VI-A Heuristic Policy with General Flow
Here we present our heuristic policy based on close-loop flow prediction (Algorithm 1). The fundamental idea is to treat the cluster of scene points near the gripper as a miniature rigid body and forecast its future movements. Then we can derive a corresponding transformation for robot execution.
@@ VI-A Heuristic Policy with General Flow
We use a RealSense D455 RGBD camera positioned behind the Franka-Emika Arm to capture an ego-view stream. The robot’s static base during manipulation acts as a reference for the FastSAM model [96] to segment the robot. More prompt points or customized models [32] can also be employed. Post-segmentation, we reconstruct 3D scene point clouds and select query points near the gripper. These points, together with the scene point clouds and the text instruction, are fed into our ScaleFlow-B model to predict the general flow. The SVD algorithm [6] is used to align the robot arm’s movement with the predicted flow. The Deoxys library [98] serves as an impedance controller for operation space control.
| Object | Action | Success Rate |
| Mug | pickup | 10 / 10 |
| putdown | 9 / 10 | |
| Toy Car | pickup | 10 / 10 |
| putdown | 10 / 10 | |
| push | 5 / 10 | |
| Clothes | fold | 8 / 10 |
| Safe | open | 9 / 10 |
| close | 10 / 10 | |
| Box | open | 10 / 10 |
| close | 10 / 10 | |
| Prismatic Furniture (Drawer) | open (pull) | 4 / 10 |
| open (grasp) | 3 / 10 | |
| close | 10 / 10 | |
| Revolute Furniture (Refrigerator) | open (pull) | 7 / 10 |
| open (grasp) | 9 / 10 | |
| close | 10 / 10 | |
| Laptop | open | 5 / 10 |
| close | 7 / 10 | |
| Average Success Rate | 81% (146 / 180) | |
VI-B Real World Experiment Setting
@@
For real-world experiments (Figure 6 and Appendix -G), we select 8 objects (covering rigid, articulated and soft bodies) across 6 scenes, encompassing 18 manipulation tasks. The rigid category includes Mug and Toy Car. Articulated objects are Safe, Box (which can be approximated as an atypical design of Safe), Laptop, Refrigerator, and Drawer, while the soft category includes Clothes. We perform “pickup” and “putdown” actions for rigid objects, with an additional “push” action for the Toy Car. Articulated objects undergo “open” and “close” tasks, and the soft object is subjected to “fold” action. (See Figure 9 in Appendix -G for visualization).
@@
As general flow affordance guides post-grasp motion, we manually position the robotic arm for task initiation. This can be replaced with automatic methods, as demonstrated in Ko et al. [50]. For storage furniture with handles, we evaluate the performance of both opened and closed grippers. Each task undergoes 10 trials, and the success rates are recorded. Discussions about the quantitative success criteria of each task can be found in Appendix -G. We also discuss the real-world baseline model [4] in Appendix -H.
VI-C Results and Analysis
@@
Transfer Ability: Does general flow facilitate stable zero-shot human-to-robot skill transfer?
Segmentation Error: How robust is the system against segmentation errors and robot occupancy?
Novel Category: Can this model generalize to the shapes of new categories that are significantly different from the training instances?
Grasp Manner: Is this object-centric system robust to variations in grasp state and gripper position?
Diverse Setting: How well can general flow adapt to diverse scenes and manipulation directions?
Augmentation Effectiveness: What impact do the designed technical augmentations have on enhancing the system’s zero-shot capabilities?
@@
Stable Zero-Shot Skill Transfer Our results, presented in Table II, demonstrate that using a general flow as a bridge enables our framework to achieve stable zero-shot human-to-robot skill transfer. An impressive 81% success rate in such challenging settings underscores the strong transfer ability of general flow in cross-embodiment robot learning. To our knowledge, this is the first flow-based work to reach such a level of zero-shot transfer performance in real-world experiments. For tasks with success rates below 60%, we meticulously analyze the reasons and propose feasible future solutions in Appendix -L.
@@
Robustness to Segmentation Error Our findings reveal that random hand mask augmentation during training significantly enhances the model’s robustness to errors in the segmentation maps of FastSAM [96]. Figure 6(a) illustrates this advantage with two examples. Notably, even with almost failed robot segmentation (as in the “open safe” task), our method still predicts meaningful flow to facilitate task completion in a closed-loop manner. Coupled with the with-hand prediction evaluation results (Table I), this evidences the strong adaptability of general flow to both human hand and incomplete Franka-Emika robot body occupancy. These findings lay the groundwork for extending general flow to a broader spectrum of robot and human data.
@@
Generalization to Novel Categories To probe the boundary of general flow’s generalization capabilities, we experiment with a “Box” category, which can be approximated as an atypical design of “Safe”, referred to “Box (Atypical Safe)”. For comparison, we also test a conventional “Safe”. Figure 6(b) presents these instances. Surprisingly, the success rate for manipulating “Box” is even higher than that of the ordinary one (100% vs. 90% for the “open” task, Table II), attributed to the structure of “Box” allowing more trajectory deviation without losing grip on the door. This underscores the strong generalization capacity of general flow methods, likely due to our model’s training on real-world datasets, which avoids any simulation-to-reality domain-gap and focuses the model on the geometric and physics features of point clouds.
@@
Robustness to Grasp Position and Manner As general flow is an embodiment-agnostic and object-centric method, it is expected to be resilient to variations in gripper position and grasp manner. To test this, we conduct manipulations using two storage pieces (a Refrigerator and a Drawer), leveraging their handles for different grasp methods. Figure 6(c) displays these different execution manners. Our model successfully completes tasks regardless of the gripper’s state. The impact of the gripper being open or closed varies with the task; for instance, a closed gripper yields a higher success rate in “open Refrigerator”. The closed gripper imposes constraints that prevent it from slipping off the door, and as a result, the impedance controller is then able to correct minor deviations in the action. Conversely, for “open Drawer”, performance decreases due to the gripper slipping on the soft leather handle. We also ensure a diverse range of gripper positions, observing minimal impact on prediction and execution efficacy. The object-centric nature of general flow lays a solid foundation for leveraging diverse policy behaviors during execution.
@@
Handling Diverse Scenes and Directions We examine the extent to which general flow can handle changes in scene and direction. Diverging from previous works’ settings [22] (demonstrated in Appendix -G), we distribute our tasks across six diverse scenes and perform scene-based prediction, eliminating the need for clean segmentation of the manipulated object. Despite these challenging conditions, our model proves robust against interference from environmentally irrelevant items. We also vary the direction of movable objects during our experiments. Figure 6(d) showcases the most challenging example in this regard. We find that our heuristic policy successfully pushes a Toy Car in different directions to a certain extent. With more data training, we anticipate the model will achieve a deeper understanding of 3D visual mechanisms and semantics.
| open Safe | close Drawer | Avg | |
| full | 9 / 10 | 10 / 10 | 95% |
| w/o Scale Rebalance | 7 / 10 | 8 / 10 | 75% |
| w/o HM Augmentation | 8 / 10 | 6 / 10 | 70% |
| w/o QPS Augmentation | 5 / 10 | 4 / 10 | 45% |
Role of Technical Augmentation Finally, we investigate the role of designed augmentations in enhancing the robustness of zero-shot transfer. We select two representative tasks (“open Safe” and “close Drawer”) for evaluation. The ablation comparisons included:
The results in Table III indicate that each augmentation significantly contributes to robust zero-shot execution. Notably, hand augmentation markedly affects tasks with substantial embodiment occupancy and occlusion, such as “close Drawer”. Query point sampling augmentation emerges as particularly influential. Currently, the PointNeXt [72] architecture inherently couples the extraction of query point features, leading to a reliance on query point sampling augmentation in our framework. We anticipate that future advancements in disentanglement architecture within 3D learning will address this issue thoroughly.
VII Conclusion
@@
In this paper, we introduce General Flow as a Foundation Affordance for scalable robot learning. For the first time, we develop a flow prediction model directly from large-scale RGBD human video datasets and successfully deploy it with a heuristic policy for stable zero-shot human-to-robot skill transfer. Our framework marks a stride in achieving scalability, universality, and stable skill transfer concurrently. We believe our work paves the way for innovative research in scalable general robot learning. In the future, we plan to further extend general flow learning to RGB datasets [16, 33], utilize depth estimation techniques [10, 50], and leverage larger RGBD datasets [34] that emerged concurrently with our work.
@@ Acknowledgments
This work is supported by the Ministry of Science and Technology of the People´s Republic of China, the 2030 Innovation Megaprojects ”Program on New Generation Artificial Intelligence” (Grant No. 2021AAA0150000). This work is also supported by the National Key R&D Program of China (2022ZD0161700). We would also like to thank Xianfan Gu for participating in the discussions of this project.
B.1 From 3D Annotated Datasets
@@
B.2 From RGBD Human Videos
Given the constraints of current RGBD Human-Object Interaction (HOI) datasets, which are either small in scale [28] or lack semantic richness [27] (mainly limited to pick & place actions), and considering the notable scarcity of resources for soft objects, we opt to collect our own RGBD videos. Coincidentally, Grauman et al. [34] releases a large-scale, semantically rich 4D HOI dataset during the same period as our project, representing a promising resource for future work. Our collection of RGBD videos for the ”fold Clothes” task, captured with a D455 depth camera, includes 30 rollouts of 6 different types of clothing, resulting in 605 extracted clips. We plan to release this dataset in the future.
@@ B.2 From RGBD Human Videos
-C Model Architecture
@@
ScaleFlow: While the main body of our paper covers the bulk of our design, we offer further details in this section. The alignment width for CLIP [73] text features is set at 6, aligning with the dimension of the original point cloud features (RGB+XYZ). In the conditional Variational Autoencoder (VAE) [49] segment, we utilize a 2-layer Multilayer Perceptron (MLP) to encode the latent variable. This is followed by another 2-layer MLP that functions as the VAE decoder. We employ separate 2-layer MLPs for scale and normalized-trajectory prediction, each featuring a hidden dimension of 512. For ScaleFlow-B, our backbone configuration mirrors that of PointNeXt-B [72]. In ScaleFlow-L, we increase the backbone width from 32 to 64. Conversely, for ScaleFlow-S, PointNeXt-B is substituted with PointNeXt-S, and the CVAE utilizes a simpler 1-layer encoder and decoder, each with a hidden dimension of 384. The loss function parameters are configured as and , with a focus on enhancing scale prediction. We maintain the latent variable dimension at 16. For more detailed information, please consult the configuration files in our code repository.
Baseline: We preserve the architecture from the original repository, with the sole modification of incorporating text features before the prediction MLP to transition the model into a multimodal version. We derive aligned text features from the original CLIP features. The dimension of these text features is set to 32. For ResNet [39] and Vision Transformer [18], we utilize the standard ‘ResNet18’ and ‘VIT-B-224’ versions, respectively. The default pretrained weights are loaded using the Timm library [87]. For R3M [65], its ‘ResNet18’ version is employed. In our architecture, all MLPs dedicated to the final flow prediction consist of 2 layers, featuring hidden dimensions of 512 and 256, respectively.
-D Training Details
@@
This section outlines additional training details not covered in the main text. For each data point, we start by calculating the center of the initial points of the general flow. We then create a cubic space centered on this point, with each side measuring 80 cm (a 40 cm perception range suffices for most tasks). To standardize point numbers for batch training, we use sampling (with the Furthest Points Sampling Algorithm) or re-sampling to obtain 2048 scene points. Standard augmentation techniques for point cloud prediction [72], including random rotation, shifting, scaling, coordinate normalization, color jittering, and feature dropping, are applied.
@@
-E Flow Visualization
We showcase the capabilities of our model, ScaleFlow-B, through visualizations of its predictions. These demonstrations include applications on original human videos as well as zero-shot real-world executions, as depicted in Figure 8. The visualizations highlight our model’s proficiency in predicting semantically and geometrically meaningful flows, even amidst the challenges posed by noisy real-world point clouds and significant embodiment occlusions. For a more comprehensive view, including zero-shot predictions for all 18 tasks, please refer to project website.
-F Quantitative Ablation Study
Experiment Setting: We conduct an ablation study to evaluate the key design elements of our methods. The variants tested include:

### v2

@@ 4 Zero-shot Real World Manipulation
In this section, our goal is to demonstrate the efficacy of general flow as an actionable guidance [29] for downstream robot manipulation. We choose one of the most challenging scenarios: zero-shot human-to-robot skill transfer in the real world. Utilizing only one prediction model for all tasks paired with a straightforward heuristic policy derived from closed-loop flow prediction, we achieve an impressive 81% average success rate. This success spans categories including rigid, articulated, and soft objects, and covers 18 tasks across 6 distinct scenes.
4.1 Heuristic Policy with General Flow
@@
Here we present our heuristic policy based on close-loop flow prediction (Algorithm 1 in Appendix). We use a RealSense D455 RGBD camera positioned behind the Franka-Emika Arm to capture an ego-view stream. The robot’s static base during manipulation acts as a reference for the FastSAM model [17] to segment the robot. More prompt points could also be employed via a designed GUI. Post-segmentation, we reconstruct 3D scene point clouds and select query points near the gripper. These query points, treated as a miniature rigid body, together with the scene point clouds and the text instruction, are fed into our to predict the general flow.
The next step involves solving the corresponding transformation, an ICP (Iterative Closest Point) problem, based on the general flow prediction. We employ the SVD (Singular Value Decomposition) algorithm [30] to align the robot arm’s end-effector transformation with the predicted flow. The choice of SVD is due to its robustness in handling noise and outliers, as well as its computational efficiency. Once the aligned transformation is computed, we use the Deoxys library [31] as an impedance controller in the operation space to execute the transformation accurately. We leave more details of deployment system and policy derivation in Appendix G.3 and Appendix G.4
4.2 Real World Experiment Setting
@@
For real-world experiments (Figure 5 and Appendix G.1), we select 8 objects (covering rigid, articulated and soft bodies) across 6 scenes, encompassing 18 manipulation tasks. The rigid category includes Mug and Toy Car. Articulated objects are Safe, Box (which can be approximated as an atypical design of Safe), Laptop, Refrigerator, and Drawer, while the soft category includes Clothes. We perform “pickup” and “putdown” actions for rigid objects, with an additional “push” action for the Toy Car. Articulated objects undergo “open” and “close” tasks, and the soft object is subjected to “fold” action. (See Figure 10, 11 in Appendix G.1 for visualization). All object categories (except for Box) and manipulation tasks appear in the training data. Except for the ”Toy Car” and ”Safe”, all other objects are novel instances for model training. To demonstrate the generalization ability of our model, we randomly build the environment backgrounds, ensuring that all evaluation settings are unseen during training.
@@
As general flow affordance guides post-grasp motion, we manually position the robotic arm for task initiation. This can be replaced with automatic methods, as demonstrated in Ko et al. [32]. For storage furniture with handles, we evaluate the performance of both opened and closed grippers. Each task undergoes 10 trials, and the success rates are recorded. Discussions about the quantitative success criteria of each task can be found in Appendix G.1. We also discuss the real-world baseline model [11] in Appendix G.2.
4.3 Results and Analysis
| Object | Action-1 | SR-1 | Action-2 | SR-2 | Action-3 | SR-3 |
| Mug | pickup | 10/10 | putdown | 9/10 | - | - |
| Toy Car | pickup | 10/10 | putdown | 10/10 | push | 5/10 |
| Clothes | fold | 8/10 | - | - | - | - |
| Safe | open | 9/10 | close | 10/10 | - | - |
| Box | open | 10/10 | close | 10/10 | - | - |
| Drawer | open (pull) | 4/10 | open (grasp) | 3/10 | close | 10/10 |
| Refrigerator | open (pull) | 7/10 | open (grasp) | 9/10 | close | 10/10 |
| Laptop | open | 5/10 | close | 7/10 | - | - |
| Average Success Rate | 81% (146 / 180) | |||||
@@ 4.3 Results and Analysis
Transfer Ability: Does general flow facilitate stable zero-shot human-to-robot transfer?
Segmentation Error: How robust is the system against robot segmentation errors?
Novel Shape: Can this model generalize to the shapes of new categories that are significantly different from the training instances?
Grasp Manner: Is this object-centric system robust to variations in grasp manner?
Diverse Setting: How well can general flow adapt to diverse scenes and setting?
@@
Stable Zero-Shot Skill Transfer Our results, presented in Table 2, demonstrate that using general flow as a bridge enables our framework to achieve stable zero-shot human-to-robot skill transfer. An impressive 81% success rate in such challenging settings underscores the strong transfer ability of general flow in cross-embodiment robot learning. To our knowledge, this is the first flow-based work to reach such a level of zero-shot transfer performance in real-world experiments. For tasks with success rates below 60%, we meticulously analyze the reasons and propose feasible future solutions in Appendix G.6. We also discuss the inference latency of our system in Appendix G.5.
@@
Robustness to Segmentation Error Our findings reveal that random hand mask augmentation during training significantly enhances the model’s robustness to errors in the segmentation maps of FastSAM [17]. Figure 5(a) illustrates this advantage with two examples. Notably, even with almost failed robot segmentation (as in the “open safe” task), our method still predicts meaningful flow to facilitate task completion in a closed-loop manner.
@@
Generalization to Novel Shape To probe the boundary of general flow’s generalization capabilities, we experiment with a “Box” category, which can be approximated as an atypical design of “Safe”. For comparison, we also test a conventional “Safe”. Figure 5(b) presents these instances. Surprisingly, the success rate for manipulating “Box” is even higher than that of the ordinary one (100% vs. 90% for the “open” task, Table 2), attributed to the structure of “Box” allowing more trajectory deviation without losing gripper on the door. This underscores the strong generalization capacity of general flow methods.
@@
Robustness to Grasp Position and Manner As general flow is an embodiment-agnostic and object-centric method, it is expected to be resilient to variations in gripper position and grasp manner. To test this, we conduct manipulations using two storage pieces (a Refrigerator and a Drawer), leveraging their handles for different grasp methods. Figure 5(c) displays these different execution manners. Our model successfully completes tasks regardless of the gripper’s state.
@@
Handling Diverse Scenes and Directions We examine the extent to which general flow can handle changes in scene and direction. We distribute our tasks across six diverse scenes and perform scene-based prediction, eliminating the need for clean segmentation of the manipulated object. We also vary the direction of movable objects during our experiments. Figure 5(d) showcases the most challenging example in this regard. We find that our heuristic policy successfully pushes a Toy Car in different directions to a certain extent.
5 Related Work
Embodiment-Agnostic Framework for Scalable Robot Learning We discuss more related works of real-world general robot learning in Appendix H. To leverage large-scale, cross-embodiment data resources [33, 34, 35, 15, 36], multiple embodiment-agnostic frameworks [37] are proposed for robot learning. Prior works [25, 38, 39, 40] employ large-scale visual pre-training to develop embodied-aware pretrained representations, but these demonstrate limited generalization [41, 42, 43]. Alternative approaches seek to derive action signals from video generation [44, 45, 46, 47, 48, 32, 49]. Affordances extracted from simulators [10, 50, 51, 52, 53] are another focus, yet they struggle with the significant sim-to-real domain-gap, particularly in 3D environments. Recent efforts [54, 11, 55, 56] attempt to directly acquire geometric-aware structured information in human video but require in-lab training or suffer from unstable performance. Instead, we leverage 3D flow-based affordance to achieve reliable zero-shot solutions.
Keypoints and Flow for Robot Learning Systems Previous studies use flow as action descriptors for robot learning [57, 58, 59, 60, 61]. However, these approaches either depend on embodiment-specific data, limiting their scalability [9, 7], or are simulation-based, facing significant sim-to-real domain-gap due to imperfect real-world RGBD point cloud generation [6, 8]. Wen et al. [62], Bharadhwaj et al. [63] utilizes flow as a prediction target for pretraining, but only operates in 2D flow and requires in-domain finetuning. In this paper, we extend the work of Seita et al. [7] into a more general version both in data resources and downstream applications. We acquire 3D flow prediction directly from RGBD human video datasets and achieve stable zero-shot skill transfer.
6 Conclusion
@@
In this paper, we introduce General Flow as Foundation Affordance for scalable robot learning. For the first time, we develop a flow prediction model directly from large-scale RGBD human video datasets and successfully deploy it with a heuristic policy for stable zero-shot human-to-robot skill transfer. Our framework marks a stride in achieving scalability, wide application, and stable skill transfer concurrently. We believe our work paves the way for innovative research in scalable general robot learning. We comprehensively discuss the limitations of our framework in Appendix I.
@@ Acknowledgments
This work is supported by the Ministry of Science and Technology of the People´s Republic of China, the 2030 Innovation Megaprojects ”Program on New Generation Artificial Intelligence” (Grant No. 2021AAA0150000). This work is also supported by the National Key R&D Program of China (2022ZD0161700).
Appendix
Appendix A Overview of Appendix
In this appendix, we offer additional implementation details and discussions of general flow. The label pipeline, code and model weights will be released in the future. Readers are also welcome to check out more details with them. This appendix is structured as follows:
Label Extraction: we delve into the pipeline specifics utilized for extracting general flow labels from RGBD human video datasets in Appendix B.
Baselines: we provide more details of our baselines in Appendix D
Ablation Study: A comprehensive quantitative ablation study is conducted in Appendix F, rigorously testing the effectiveness of our algorithmic design.
Real-World Experiment: This section is dedicated to an expansive elucidation of real-world experiments, encompassing the experimental setup (Appendix G.1), real-world baseline comparisons (Appendix G.2), robot system development (Appendix G.3), policy derivation strategies (Appendix G.4), inference latency measurements (Appendix G.5), and analysis of failure cases (Appendix G.6).
Related Work of Real-world General Robot Learning: we further discuss the related works of general robot policy training in Appendix H.
Limitations: We comprehensively discuss the limitations of our current framework in Appendix I, including data diversity, policy backbone, manual grasping, representation format and heuristic policy.
Codebases: Acknowledgements are extended to the multiple codebases that have been instrumental in supporting this project in Appendix J.
Additional videos and flow visualizations are available https://general-flow.github.io.
Appendix B Label Extraction Pipeline
General flow labels can be directly extracted from 3D human datasets or RGBD videos. Figure B.2 displays some data resources we utilize.
B.1 From HOI4D Datasets
@@
B.2 From Collected RGBD Videos
Given the constraints of current RGBD Human-Object Interaction (HOI) datasets, which are either small in scale [34] or lack semantic richness [36] (mainly limited to pick & place actions), and considering the notable scarcity of resources for soft objects, we opt to collect our own RGBD videos. Our collection of RGBD videos for the ”fold Clothes” task, captured with a D455 depth camera, includes 30 rollouts of 6 different types of clothing, resulting in 605 extracted clips. We plan to release this dataset in the future.
@@ B.2 From Collected RGBD Videos
Appendix C Model Architecture and Training
C.1 Architecture
@@
While the main body of our paper covers the bulk of our design, we offer further details in this section. The alignment width for CLIP [19] text features is set at 6, aligning with the dimension of the original point cloud features (RGB+XYZ). In the conditional Variational Autoencoder (VAE) [21] segment, we utilize a 2-layer Multilayer Perceptron (MLP) to encode the latent variable. This is followed by another 2-layer MLP that functions as the VAE decoder. We employ separate 2-layer MLPs for scale and normalized-trajectory prediction, each featuring a hidden dimension of 512. For ScaleFlow-B, our backbone configuration mirrors that of PointNeXt-B [20]. In ScaleFlow-L, we increase the backbone width from 32 to 64. Conversely, for ScaleFlow-S, PointNeXt-B is substituted with PointNeXt-S, and the CVAE utilizes a simpler 1-layer encoder and decoder, each with a hidden dimension of 384. The loss function parameters are configured as and , with a focus on enhancing scale prediction. We maintain the latent variable dimension at 16. For more detailed information, please consult the configuration files in our code repository.
C.2 Augmentation Techniques
In light of the complex environmental challenges encountered in zero-shot real-world deployments, we propose two technical augmentations to boost zero-shot generalization robustness. The fundamental concept involves simulating various states of hand occlusion and distributions of query points, compelling the model to adapt to diverse conditions:
Hand Mask (HM) Augmentation: We encounter occlusions from human hands in our training data while facing occlusions from robot arms during deployments. Therefore, it is crucial to enhance the model’s resilience to embodiment occlusions. To achieve this, we manipulate the presence of points on the hand in the input scene point clouds. We choose one of three rules, with probabilities =0.5, =0.2, and =0.3: (1) deleting all hand points; (2) keeping all hand points; and (3) sampling a random anchor point on the hand and retaining only points that are more than 12cm away from the anchor. These three rules are designed to simulate situations where the gripper experiences no occlusion, full occlusion, and partial occlusion during robot execution.
Query Points Sampling (QPS) Augmentation: Different downstream applications may require varying query point sampling methods. Consequently, our model must be adaptable to various query point distributions. We achieve this by augmenting the training process. In each training iteration, we select a subset of available query points using one of two rules, based on probabilities =0.7, =0.3: (1) complete random sampling; (2) randomly selecting an anchor query point and then choosing a specific number of points closest to the anchor. The complete random sampling ensures that all points are covered, whereas anchor-based sampling maintains a structure more aligned with the format used in downstream action guidance generation.
C.3 Training Details
@@
We utilize 1.5s video clips as our training data and set the time steps of general flow to 3 for all data sources. The dataset is divided into training, validation, and test sets in an 80%, 10%, 10% ratio, resulting in 51693, 6950, and 6835 clips respectively, with no identical object instances across sets. Each sample consists of 2048 scene points sampled in an 808080 cube space around the center of the flow start points using the furthest point sampling (FPS) algorithm (a 40 cm perception range suffices for most tasks). During training, we randomly sample 128 query points, while for validation, 512 points are randomly sampled. Standard augmentation techniques for point cloud prediction [20], including random rotation, shifting, scaling, coordinate normalization, color jittering, and feature dropping, are applied.
@@
Given that our datasets include ground-truth labels for object parts, we distribute 512 query points across each part equally during testing to enhance evaluation effectiveness. It’s important to note that we do not use any part labels in model validation and real-word testing.
Appendix D Baselines
We adapt three types of relevant work to our setting:
We preserve the architecture from the original repository, with the sole modification of incorporating text features before the prediction MLP to transition the model into a multimodal version. We derive aligned text features from the original CLIP features. The dimension of these text features is set to 32. For ResNet [23] and Vision Transformer [24], we utilize the standard ‘ResNet18’ and ‘VIT-B-224’ versions, respectively. The default pretrained weights are loaded using the Timm library [65]. For R3M [25], its ‘ResNet18’ version is employed. In our architecture, all MLPs dedicated to the final flow prediction consist of 2 layers, featuring hidden dimensions of 512 and 256, respectively. To ensure a fair comparison, scale rebalance, HMA augmentation and QPS augmentation are applied to all baselines.
Appendix E Emergent Properties of General Flow
| Model | ADE | FDE | ADE-H | FDE-H | Param (M) |
| ScaleFlow-S | 3.74 | 5.01 | 3.72 | 4.98 | 0.906 |
| ScaleFlow-B | 3.58 | 4.77 | 3.56 | 4.74 | 5.622 |
| ScaleFlow-L | 3.55 | 4.70 | 3.52 | 4.67 | 17.088 |
E.1 Properties Analysis
In this section, we demonstrate the notable emergent properties of general flow such as robustness to hand occlusion, language-driven semantic control, resilience to label noise, and scale spatial adaption.
Robustness to hand occlusion We first test robustness to hand occupancy in inputs across all 3D models, which are denoted as ADE-H and FDE-H (evaluations that include hand points in the model’s input). As shown in Table 3, with appropriate augmentations, our model are robust for hand occupancy.
language-driven semantic control Through large-scale training, our model not onlyaptures rich semantic information but also becomes adeptly controllable through language modality. As depicted in Figure 8(a)(b), our model demonstrates the capability to predict varied flows for identical input point clouds when provided with different instructions.
Resilience to label noise Furthermore, it is remarkably robust to label noise. Figure 8(c)(d) showcases two instances of this resilience: despite severe label noise (notable deviation in “open Safe” and near-static in “pickup Toy Car”), our model accurately predicts the correct trend.
Scale spatial adaption Additionally, our model could perform scale spatial adaption through scalable training. It dynamically adjusts its prediction scale in response to the spatial relationships of objects, such as ending on the table and scaling up for longer distances, as seen in Figure 8(e)(f). All these emerging phenomena reveal the benefits of large-scale training.
E.2 Real-Robot Validation
We also validate the prediction properties through real-robot experiments. we confirm the ”scale spatial adaptation” phenomenon in the ”putdown Mug” task in a real-robot setting, as depicted to the right of Figure 9.
We then demonstrate the model’s ability for language-controlled multi-semantic action generation. We select ”Drawer,” ”Refrigerator,” and ”Mug” as manipulation objects, setting their initial states the same while varying only the instructional verbs for task execution. These results, presented to the left of Figure 9, illustrate that language instructions enable the execution of various behaviors within a single scene.
Appendix F Ablation Study
F.1 Dataset Experiment Setting
We conduct an ablation study to evaluate the key design elements of our methods. The variants tested include: